# A Data-Centric Perspective on Evaluating Machine Learning Models for Tabular Data

### v2

@@ D.1 Comparison of Preprocessing Pipelines per Model and Dataset
D.2 Analysis of Modeling Components
@@
These results are in line with common practice in ML competitions [81]. While recent work strongly focuses on model selection [35, 63], participants of Kaggle competitions typically stick to few model classes and instead focus on developing feature engineering techniques for these particular models [42]. Predictive Machine Learning is a winner-takes-all game. Selecting another model than the one that is known to work best is only important if the default model fails or if ensembling is necessary. Hence, it makes sense to look at the problem from the winner-takes-all perspective and to evaluate performance gains from different modeling decisions w.r.t. a strong baseline. It is known from related work that CatBoost is the strongest model with default hyperparameters [63]. Hence, we evaluate performance gains over a CatBoost baseline. Figure 9 illustrates average leaderboard position gains over the baseline for different modeling decisions. It can be seen that without feature engineering, the best average leaderboard position is the 14.5% percentile, while without model selection, it is the 3% percentile. Hence, one of the most important takeaways is that current tabular ML research overemphasizes model evaluation but underestimates data preprocessing especially feature engineering. While TTA increases the average leaderboard position by 8.3% in the default setting, its importance after model selection and hyperparameter optimization reduces. Without TTA, the best achievable average leaderboard position was 3.2% and 1.6% with TTA, indicating a relatively small but important effect. In this Figure, the average top position is the 1.6% percentile because we only considered single models. The missing component for scoring in the top 1% percentile is ensembling, which we have shown to achieve using AutoGluon in the main paper.
@@
D.3 Using Our Framework for Evaluating New Methods
@@
In Section 5 we identified directions for future work. These directions were based on general insights of our analysis for the tabular data field. This subsection will showcase how our framework can be utilized to develop new approaches and compare them to expert solutions. Our framework contains, to our knowledge, the largest collection of implemented expert solutions for relevant datasets. Hence, our framework is especially useful to researchers developing AutoML solutions, especially focusing on feature engineering. Furthermore, our framework can be useful to researchers developing model-specific and data-agnostic preprocessing pipelines, i.e., for novel neural networks. In addition, our framework can be used to develop test-time adaptation methods for tabular data. Figure 10 shows four particular challenges for future work in tabular Deep Learning and AutoML for which our framework can serve as a benchmark to measure progress: A) Develop a neural network not relying on feature engineering techniques; B) Develop a neural network capable of test-time adaptation to replace the often infeasible test-time feature engineering. C) Create a universal model-agnostic automated feature engineering pipeline that surpasses expert feature engineering; D) Enhance AutoML solutions to outperform expert feature engineering pipelines;
@@
Figure 10: Challenges for further automating deep learning and AutoML for tabular data. Challenge A compares the best neural networks within the standardized and feature engineering pipelines after extensive HPO. Challenge B compares the best neural networks within the feature engineering and the test-time adaptation pipeline after extensive HPO. Challenge C compares the best model within the standardized pipeline to the best model within the feature engineering pipeline. Challenge D compares AutoGluon to the best model within the feature engineering pipeline.
Appendix E Limitations
@@
  • We use Kaggle competitions in an effort to evaluate more realistic tasks than evaluated in related work. It is important to highlight that the competition setup on Kaggle does not always reflect real-world tasks. However, due to the involvement of companies and institutions and the poor availability of high-quality tabular datasets [47], these are arguably among the most realistic datasets available as open-source data. Furthermore, one of our contributions was to separate aspects from the main learning task that made competitions unrealistic (i.e., data leaks or, for some applications, test-time adaptation). This additionally improves the real-world transferability of our experiments.
@@
  • Our evaluation framework does not allow to assess whether one model is generally better than another model. We only claim that model comparisons change and that feature engineering and preprocessing greatly influence model comparison on our datasets. For a more generalizable model comparison using more datasets, we refer to related work [35, 63].
  • Due to the extent of our experiments (over 200,000 trained models), it was infeasible to repeat the experiments multiple times to obtain error bars. Nevertheless, our experiments include randomness (e.g., CV splits, weight initialization for deep learning models, or bagging for the tree-based models), limiting the generalizability of our results. However, the extent of our experiments and the clear differences between the implemented preprocessing pipelines over multiple models and datasets make the risk of randomness affecting our claims very low despite not being explicitly quantified for all models.
@@ Appendix F Evaluation on the Original Task Metrics
This Section lists the main results using the original task metrics for all experiments. An overview of important components per dataset and model can be seen in Figure 11. All experimental results on the original metrics can be seen in Tables 18, 19, 20, and 21. Further results can be seen in our code. Overall, the evaluation using the original metrics aligns with our main findings.

### v3

@@ D.3 Comparison of Preprocessing Pipelines per Model and Dataset
D.4 Analysis of Modeling Components
@@
These results are in line with common practice in ML competitions [68]. While recent work strongly focuses on model selection [26, 55], participants of Kaggle competitions typically stick to few model classes and instead focus on developing feature engineering techniques for these particular models [35]. Predictive Machine Learning is a winner-takes-all game. Selecting another model than the one that is known to work best is only important if the default model fails or if ensembling is necessary. Hence, it makes sense to look at the problem from the winner-takes-all perspective and to evaluate performance gains from different modeling decisions w.r.t. a strong baseline. It is known from related work that CatBoost is the strongest model with default hyperparameters [55]. Hence, we evaluate performance gains over a CatBoost baseline. Figure 10 illustrates average leaderboard position gains over the baseline for different modeling decisions. It can be seen that without feature engineering, the best average leaderboard position is the 14.5% percentile, while without model selection, it is the 3% percentile. Hence, one of the most important takeaways is that current tabular ML research overemphasizes model evaluation but underestimates data preprocessing especially feature engineering. While TTA increases the average leaderboard position by 8.3% in the default setting, its importance after model selection and hyperparameter optimization reduces. Without TTA, the best achievable average leaderboard position was 3.2% and 1.6% with TTA, indicating a relatively small but important effect. In this Figure, the average top position is the 1.6% percentile because we only considered single models. The missing component for scoring in the top 1% percentile is ensembling, which we achieve using AutoGluon in the main paper.
@@
D.5 Evaluation At Different Snapshots of the Competitions
In the main paper, we evaluated results w.r.t. the end of the competition as a reference snapshot. All publicly available metadata from Kaggle competitions is available at Meta Kaggle,(https://www.kaggle.com/datasets/kaggle/meta-kaggle). This allows us to additionally use different points in time when the competition took place. Table 19 shows the performance of the best model in the standardized preprocessing pipeline after varying number of days in the competition. For most datasets, scoring high positions at the first day without feature engineering is possible with our framework. This shows that our framework can be of use for participants in Kaggle competitions to obtain first baseline results.
1 7 14 21 28 45 60 end
MBGM Best single 1.0 0.998 0.997 0.996 0.997 0.994 0.994 0.994
AutoGluon 0.87 0.814 0.706 0.67 0.653 0.616 0.616 0.616
SVPC Best single 1.0 1.0 1.0 1.0 1.0 0.991 0.964 0.946
AutoGluon 1.0 1.0 1.0 1.0 1.0 0.985 0.873 0.794
AEAC Best single 1.0 1.0 0.996 0.991 0.989 0.982 0.958 0.953
AutoGluon 0.98 0.894 0.781 0.739 0.698 0.667 0.638 0.616
OGPCC Best single 0.996 0.994 0.987 0.984 0.977 0.963 0.908 0.896
AutoGluon 1.0 1.0 1.0 0.999 1.0 0.998 0.996 0.996
SCS Best single 1.0 0.978 0.972 0.95 0.905 0.865 0.825 0.825
AutoGluon 1.0 0.972 0.966 0.936 0.885 0.837 0.792 0.792
BPCCM Best single 1.0 0.998 0.996 0.997 0.997 0.993 0.991 0.986
AutoGluon 1.0 1.0 0.998 0.997 0.997 0.997 0.995 0.992
SCTP Best single 0.991 0.895 0.893 0.865 0.681 0.551 0.511 0.511
AutoGluon 0.986 0.867 0.865 0.822 0.655 0.531 0.493 0.493
HQC Best single 1.0 1.0 1.0 1.0 0.998 0.999 0.995 0.992
AutoGluon 1.0 1.0 1.0 1.0 1.0 1.0 0.995 0.993
IFD Best single 1.0 0.991 0.985 0.978 0.922 0.776 0.694 0.621
AutoGluon 0.14 0.145 0.165 0.185 0.187 0.178 0.17 0.192
PSSDP Best single 0.995 0.969 0.889 0.845 0.781 0.696 0.638 0.638
AutoGluon 0.968 0.889 0.786 0.727 0.667 0.593 0.552 0.552
Table 19: Private leaderboard position at different points in time of the competitions. For each competition, the performance of the best model trained in the standardized preprocessing pipeline is reported. The columns represent days after the competition started.
D.6 Using Our Framework for Evaluating New Methods
@@
In Section 5 we identified directions for future work. These directions were based on general insights of our analysis for the tabular data field. This subsection will showcase how our framework can be utilized to develop new approaches and compare them to expert solutions. Our framework contains, to our knowledge, the largest collection of implemented expert solutions for relevant datasets. Hence, our framework is especially useful to researchers developing AutoML solutions, especially focusing on feature engineering. Furthermore, our framework can be useful to researchers developing model-specific and data-agnostic preprocessing pipelines, i.e., for novel neural networks. In addition, our framework can be used to develop test-time adaptation methods for tabular data. Figure 11 shows four particular challenges for future work in tabular Deep Learning and AutoML for which our framework can serve as a benchmark to measure progress: A) Develop a neural network not relying on feature engineering techniques; B) Develop a neural network capable of test-time adaptation to replace the often infeasible test-time feature engineering. C) Create a universal model-agnostic automated feature engineering pipeline that surpasses expert feature engineering; D) Enhance AutoML solutions to outperform expert modeling pipelines;
@@
Figure 11: Challenges for further automating deep learning and AutoML for tabular data. Challenge A compares the best neural networks within the standardized and feature engineering pipelines after extensive HPO. Challenge B compares the best neural networks within the feature engineering and the test-time adaptation pipeline after extensive HPO. Challenge C compares the best model within the standardized pipeline to the best model within the feature engineering pipeline. Challenge D compares AutoGluon to the best submission after 24 hours and the best models within the feature engineering pipeline.
Appendix E Discussion of Limitations
@@
  • We use Kaggle competitions in an effort to evaluate more realistic tasks than in related work. It is important to highlight that the competition setup on Kaggle does not always reflect real-world tasks. However, due to the involvement of companies and institutions and the poor availability of high-quality tabular datasets [41], these are arguably among the most realistic datasets available as open-source data. Furthermore, one of our contributions was to separate aspects from the main learning task that made competitions unrealistic (i.e., data leaks or, for some applications, test-time adaptation). This additionally improves the real-world transferability of our experiments.
@@
  • It is worth mentioning that the implemented feature engineering steps for each dataset were always extracted from modeling pipelines were they worked well in combination with specific models, mostly tree-based. It might be the case that some models would additionally benefit from other preprocessing techniques that were not part of the implemented pipeline. A general observation was that feature engineering techniques working well for one model often also work well for others. Only for one dataset, the expert solution (slightly) differentiated between tree-based and neural network preprocessing.
@@
  • Another possible issue with using the leaderboard as a metric is the suggestive wording possibly leading to misinterpretations of what expert solutions in our context are. First, not all submitted solutions are expert solutions and the leaderboard is skewed. Users of our framework should avoid formulations such as "beating 99% of experts", as it is unclear which of the submitted solutions can be considered expert level. Instead, factually correct statements such as "top 1% of all competition participants" should be used.
  • Our evaluation framework does not allow to assess whether one model is generally better than another model. We only claim that model comparisons change and that feature engineering and preprocessing greatly influence model comparison on our datasets. For a more generalizable model comparison using more datasets, we refer to related work [26].
  • Due to the extent of our experiments (over 200,000 trained models), it was infeasible to repeat the experiments multiple times to obtain error bars. Nevertheless, our experiments include randomness (e.g., CV splits, weight initialization for deep learning models, or bagging for the tree-based models), limiting the generalizability of our results. However, the extent of our experiments and the clear differences between the implemented preprocessing pipelines over multiple models and datasets make the risk of randomness affecting our main claims very low despite not being explicitly quantified for all models.
@@ Appendix F Evaluation on the Original Task Metrics
This Section lists the main results using the original task metrics for all experiments. An overview of important components per dataset and model can be seen in Figure 12. All experimental results on the original metrics can be seen in Tables 20, 21, 22, and 23. Further results can be seen in our code. Overall, the evaluation using the original metrics aligns with our main findings.